OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

OpenStudy (anonymous):

@KeithAfasCalcLover can you help?

OpenStudy (anonymous):

I might :) So you know \(y=mx+b\). Here you see that you have two points \(P_1(4,0)\) And \(P_2(2,6)\). Have you been introduced to the formula for slope?

OpenStudy (anonymous):

A.y = 4x – 2 B. y = 2x – 4 C. y = –3(x – 4) D. y = 3(x + 4)

OpenStudy (anonymous):

yes I have @KeithAfasCalcLover

OpenStudy (anonymous):

Yeah so you know that the slope, \(m\) is \[m=\frac{y_2-y_1}{x_2-x_1}\] And so we know that the slop here is: \[m=\frac{6-0}{2-4}=\frac{6}{-2}=-3\] So we know we have the equation \[y=-3x+b\] we just need to know our \(b\) which we can find out by substituting a point:

OpenStudy (anonymous):

So we can sub in the first or second, your pick ;)

OpenStudy (anonymous):

okay now what?

OpenStudy (anonymous):

Well its up to you, want to sub in the first or second point? doesn't matter really!

OpenStudy (anonymous):

1st one!